[figure: 9.7 cm]

ONE DIMENSIONAL SPIN-GLASS-LIKE BEHAVIOR OF THE

NONCOLLINEAR ISING COMPOUND CsDy(MoO_{4})_{2}?

E. Khatsko^{a} C. Paulsen^{b} and A.Cherny^{a}

aInstitute for Low Temperature Physics, 47 Lenin Ave. 61103 Kharkov, Ukraine bCRTBT, Associe a lUniversite Joseph Fourier CNRS BP 166 38042 Grenoble Cedex 9,

France

Caesium dysprosium molybdate CsDy(MoO_{4})_{2} belong to the family of low-dimensional Ising compounds with rhombic symmetry of crystal lattice. At T_{c} = 1.3K the long order magnetic phase transition in noncollinear magnetic structure take place. This transition is confirmed by neutron scattering experiments. Nevertheless, at low temperature (3.6-0.3 K)the dc susceptibility along the c-axis is largely dependent of the cooling mode: zero field cooled or field cooled. Additionally ac susceptibility and time dependent magnetisation measurement in this direction show reminiscent spin-glass behavior with anomalous long relaxation time. No anomalies are found in a- and b-directions. Possible models are discussed.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : oral

Corresponding author : E. Khatsko

Address for correspondence : Institute for Low Temperature Physics, 47 Lenin Ave. 61103 Kharkov, Ukraine

Email address : khatsko@ilt.kharkov.ua